OpenStudy (anonymous):

pllzzz show me how to get the answer What is the solution to the equation? -6.8+2.6+x=9.7 x=?

pooja195 (pooja195):

Start by combining like terms \[\huge~\rm~\bf(x)+(−6.8+2.6)=9.7\]

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Did u find the equation you are left with?

OpenStudy (anonymous):

(x)+-4.2=9.7 is that correct? @pooja195

pooja195 (pooja195):

\[\Huge~\rm~x−4.2=9.7\] now just add 4.2 to both sides x=9.7+4.2

OpenStudy (anonymous):

13.9? @pooja195

pooja195 (pooja195):

Good job :)

OpenStudy (anonymous):

thank youuu

pooja195 (pooja195):

yw ^-^